OpenStudy (anonymous):

Can someone please help me find an equation for this word problem? David and keith are route drivers for a fast photocompany. David's route is 80 miles, and keith 's is 100 miles. Keith averages 10 mph more than David and finishes his route 10 minutes before David. What is David's speed?

OpenStudy (anonymous):

one say is \[\frac{100}{r+10}=\frac{80}{r}+10\]

OpenStudy (anonymous):

*way

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

i think i made a mistake, let me check

OpenStudy (anonymous):

all these names confuse me. David's rate is r, so Keith's rate is r + 10

OpenStudy (anonymous):

david's time is therefor \[\frac{80}{r}\] and keith's time is \[\frac{100}{r+10}\]

OpenStudy (anonymous):

so yeah i screwed up the next part. Keith is ten minutes faster so it should be \[\frac{100}{r+10}=\frac{80}{r}-10\]